OpenStudy (anonymous):

Can someone help me simplify: 5/3=((14-L)^2-(12-L)^2))/((12-L)^2-(10-L)^2)

OpenStudy (anonymous):

gimmick here is that both numerator and denominator are the difference of two square

OpenStudy (anonymous):

\[a^2-b^2=(a+b)(a-b)\] so \[(14-L)^2-(12-L)^2=(14-L+12-L)(14-L-(12-L)\]

OpenStudy (anonymous):

which, combining like terms then multiplying gives \[(26-2L)\times 2=52-4L\]

OpenStudy (anonymous):

denominator is similar

OpenStudy (anonymous):

wow so simple haha thanks a lot @satellite73 its an example of finding the natural length of a spring problem totally helps.

OpenStudy (anonymous):

so for some odd reason i cant get the denominator right

OpenStudy (anonymous):

nevermind got it!